what happens to the (1) resistance and (2) current flow through a conductor if:

b. conductivity increases

c. decrease in temperature


1
Expert's answer
2022-04-11T12:05:29-0400
"R=\rho\frac{l}{A},\quad R=R_0(1+\alpha \Delta T),\quad I=\frac{V}{R}"

(a) resistance decreases, current increases.

(b) resistance decreases, current increases.